OpenStudy (anonymous):

A jar contains 8 green, 5 red, 4 blue, and 3 black marbles.once a marble is selected, it is not replaced. find each probability. 1) p(two green marbles) 2) p(two black marbles) 3) p(a red marble and then a blue marble) 4) p(a black marble then a green marble) 5) p(two marbles that are not red) 6) p(two marbles that are neither blue nor green) 7) a family has two children, both of whom are boys. what is the probability that the next child will be born a boy?

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

im new hear im pretty much lost

OpenStudy (anonymous):

pick one want to do the first one?

OpenStudy (anonymous):

1)

OpenStudy (anonymous):

pick one marble how many are green?

OpenStudy (anonymous):

i mean in total how many green ones are there?

OpenStudy (anonymous):

8 green marbles

OpenStudy (anonymous):

out of how many total?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

actually the total is 20, the total includes the green ones

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

therefore the probability you pick a green one on the fist try is \(\frac{8}{20}\) is that much okay?

OpenStudy (anonymous):

but it says they picked out 2 green marbles

OpenStudy (anonymous):

yes i know, but we have to do it one step at a time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first we pick a marble and the probability it is green is \(\frac{8}{20}\) then we pick another marble. assuming that the first one we picked was green, how many green marbles are left in the bag?

OpenStudy (anonymous):

7 out of 19

OpenStudy (anonymous):

yes, there are 7 green ones left, and 19 total in the bag, so the probability that the second one is green given that the first one was also green is \(\frac{7}{19}\) what you said

OpenStudy (anonymous):

now to compute the probability that BOTH are green, take those two numbers and multiply them

OpenStudy (anonymous):

i.e. it is \(\frac{8}{20}\times \frac{7}{19}\)

OpenStudy (anonymous):

thanks for your help but i really need to move on. i have alot of homework to do. thank you for your help. bye

OpenStudy (anonymous):

yw all the others are exactly the same idea

OpenStudy (anonymous):

i still dont get it but anyways thanks for trying